The body is thrown vertically upward at a speed of 20 m / s. At what height the kinetic energy of the body will be equal to its potential energy. Disregard air resistance


1
Expert's answer
2021-04-16T07:23:36-0400

We have a condition for a certain height, where


"PE=KE."


From conservation of energy, we have


"KE_i=PE+KE,\\
KE_i=PE+PE=2PE.\\\space\\
\frac12mv^2=2·mgh,\\\space\\
h=\frac{v^2}{4g}=10.2\text{ m}."